OpenStudy (anonymous):

hii what is the most general antiderivative of the function F(x)= x-4?

OpenStudy (anonymous):

antiderivative of a function is the integral of a function, therefore \[\int\limits_{}^{}F(x)dx= \int\limits_{}^{}( x-4)dx=(x^2)/2 -4x +C\]

OpenStudy (anonymous):

Thank you so much!!! how in the world do you know how to do that?! I HATE CALCULUS!!!

OpenStudy (anonymous):

integral is the reverse of derivative \[\int\limits_{}^{}F(x)dx=\int\limits_{}^{} (x-4)dx\]

OpenStudy (anonymous):

\[=\int\limits_{}^{}xdx-\int\limits_{}^{}4dx\] =(x^2)/2 - 4x +C

OpenStudy (anonymous):

lets do the derivative of your answer d(x^2)/2 - 4x +C)/dx=(2/2)x -4+0 =x-4=F(x)